a paddle boat can move at the speed of 20 km/h in still water. the boat is paddled 12km downstream in the same time it takes to go 6km upstream. what is the speed of the river.
Answer by mananth(16949)

a paddle boat can move at the speed of 20 km/h in still water. the boat is paddled 12km downstream in the same time it takes to go 6km upstream. what is the speed of the river.
let the speed of the river be x mph
downstream:
distance 12 kms
speed = x+20 km/hr
Time = 12/ x+20
..
Upstream
Distance = 6 km
speed = 20-x km/hr
Time = 6/ 20-x
.
Time taken is same
12/ x+20 = 6 / 20-x
12(20-x)=6(x+20)
240-12x= 6x+120
-18x=-120
x= 6.66 km/hr the speed of the river